# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v1

@@ 5 Results
5.1 Revisiting Speed, Utilization and Stalls
Using our experimental configurations from Section 1, we verify the efficacy of Algorithm 2 compared to Algorithm 1. We implement Algorithm 2 in Triton 10.1145/3315508.3329973 and run our experiments on a 4060 Ti. First, we compare the backward pass of the FlashKAT group-wise rational against the KAT group-wise rational by measuring cycles, time, and throughput utilization for SMs and memory. As shown in Table 3, the FlashKAT group-wise rational achieves a speedup of 140.5x over the KAT group-wise rational, with improvement across the utilization of SM, cache, and memory, contributing to the reduction in backward pass time.
Table 3: A comparison between the group-wise rational backward pass for KAT and FlashKAT using FLOPs, cycles, time, SM, and memory throughput.
@@ 5.1 Revisiting Speed, Utilization and Stalls
0011112222Mem Stall - Long ScoreboardComputing - Selected2.312.312.312.311111Cycles
Figure 3: The warp-state statistics for the FlashKAT group-wise rational backward pass. "Computing - Selected" is where the warp does useful computation; all the others are memory-stall states.
Secondly, we conduct again the analysis of warp state statistics for the backward pass of FlashKAT group-wise rational. From observing the warp state statistics in Figure 3, we can see that only the "Stall Long Scoreboard" is still longer than the "Selected" compute state; all the other 8 memory stalls are shorter than compute. Furthermore, the "Stall Long Scoreboard" has reduced from 981.51 cycles to 2.31 cycles. Such a result indicates that the new kernel is now spending more time performing computation rather than mostly waiting for memory transfers.
5.2 Image Recognition Training Speed and Accuracy
We validate that the training capabilities of FlashKAT are consistent with KAT by training it on ImageNet-1k for image recognition (russakovsky2015imagenet, ). While training we also measure the training throughput of FlashKAT to compare against KAT, ViT, and DeiT. The data loader time is excluded from this measurement as it is dependent on external factors (e.g., CPU, main memory, etc.). We measure the throughput in images/second as the average of 100 samples, with 95% confidence intervals provided.
Our training procedure follows an identical curriculum to yang2024kolmogorov , wherein we follow the hyperparameters of DeiT touvron2021training . As such, we train using a batch size of 1024 with the AdamW optimizer loshchilov2017decoupled . In GR-KAN, each of the 8 groups shares its 6 numerator coefficients and has 4 unique denominator coefficients. The attention layers are initialized with Mimetic initialization trockman2023mimetic . The first layer of GR-KAN has its group-wise rational initialized to the identity function, and the second layer is initialized to a Swish function ramachandran2017searching . The patch size is set to 16. For data augmentation and regularization techniques, we apply RandAugment cubuk2020randaugment , Mixup zhang2017mixup , CutMix yun2019cutmix , Random Erasing zhong2020random , weight decay, Label Smoothing szegedy2016rethinking and Stochastic Depth huang2016deep . All experiments are conducted using an H200 GPU. Appendix A.1 provides the complete set of model configurations and training hyperparameters.
@@ 5.2 Image Recognition Training Speed and Accuracy
5.3 Reduced Gradient Rounding Error
@@
In addition to a significant speedup, the proposed FlashKAT also offers reduced rounding error when computing the gradients for the learned group-wise rational coefficients. This is a byproduct of our method, accumulating gradients in each block through a reduction sum rather than using atomic adds. Table 5 demonstrates that the mean absolute error in gradient outputs in FlashKAT is orders of magnitude smaller than KAT. As such, FlashKAT is less susceptible to rounding errors that may contribute towards unstable gradient updates (e.g. low-precision training). Due to space limitations, more details are provided in Appendix A.2.
@@ 6 Conclusion
The successes of the KAN, due to its increased expressivity and interpretability, have made it a popular alternative to the MLP. However, it has yet to achieve widespread application in large-scale NLP and Vision tasks due to its increased parameters, computational cost, lack of GPU hardware optimization, and training instability. The KAT has resolved many of the KAN’s shortcomings with the introduction of GR-KAN, but it still suffers from slow training speeds. In this work, we conduct experimentation to understand the source of the slowdown and identify that the slow training speeds are a byproduct of memory stalls in the backward pass. To address this bottleneck, we propose a novel restructuring of the group-wise rational that allows efficient gradient accumulation by minimizing atomic add usage. The resulting FlashKAT achieves an 86.5x speedup compared to ViT, while reducing gradient rounding errors. The application of our work both enables wider adoption and expedites research on the KAT.
@@ Appendix A Appendix
A.1 Image Recognition Training Hyperparameters
@@
Model Layers Hidden size (D) MLP size Heads Params
@@
Adam $\epsilon$ $1\times 10^{-8}$ $1\times 10^{-8}$ $1\times 10^{-8}$
@@
Learning rate $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$
@@
A.2 Reduced Gradient Rounding Error
FlashKAT also reduces rounding error when computing the gradients for the learned group-wise rational coefficients, a byproduct of our implementation accumulating gradients in each block through sum reduction rather than using atomic adds. We validate that FlashKAT reduces gradient rounding errors by constructing an experiment measuring the coefficient gradient rounding errors. Our experiment generates $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, and $\mathbf{B}\in\mathbb{R}^{8\times 4}$ whose entries are distributed according to $\mathcal{N}(0,1)$. Then we compute $\mathbf{dA}\in\mathbb{R}^{8\times 6}$ and $\mathbf{dB}\in\mathbb{R}^{8\times 4}$ using the KAT group-wise rational backward pass method in float64 and float32, and the FlashKAT group-wise rational backward pass method in float32. We measure the mean absolute error (MAE) over 100 passes between the float64 gradient output and two float32 outputs. All these experiments were done on an RTX 4060 Ti. We report the mean, 95% confidence interval, and variance in Table 8.

### v2

@@ 5 Results
Revisiting Speed, Utilization and Stalls
Using our experimental configurations from Section 3, we verify the efficacy of Algorithm 2 compared to Algorithm 1. We implement Algorithm 2 in Triton (10.1145/3315508.3329973) and run our experiments on a 4060 Ti. First, we compare the backward pass of the FlashKAT group-wise rational function against the KAT group-wise rational function by measuring cycles, time, and throughput utilization for SMs and memory. As shown in Table 3, the FlashKAT group-wise rational function achieves a speedup of 140.5x over the KAT group-wise rational function, with improvements across the utilization of SM, cache, and memory, contributing to the reduction in backward pass time.
Table 3: A comparison between the group-wise rational function backward pass for KAT and FlashKAT using FLOPs, cycles, time, SM, and memory throughput.
@@
01122MS - Long ScoreboardComputing - Selected2.312.3111Cycles
Figure 3: The warp‐state statistics for the FlashKAT group-wise rational function backward pass. ”Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.
Secondly, we again conduct the warp state statistics analysis for the backward pass of the FlashKAT group-wise rational function. From observing the warp state statistics in Figure 3, we can see that only the ”Stall Long Scoreboard” is still longer than the ”Selected” compute state; all the other 8 memory stalls are shorter than compute. Furthermore, the ”Stall Long Scoreboard” has reduced from 981.51 cycles to 2.31 cycles. Such a result indicates that the new kernel is now spending more time performing computation rather than mostly waiting for memory transfers.
Image Recognition Training Speed and Accuracy
@@
We validate that the training capabilities of FlashKAT are consistent with KAT by training it on ImageNet-1k for image recognition (russakovsky2015imagenet). While training, we also measure the training throughput of FlashKAT to compare against KAT, ViT, and DeiT. The data loader time is excluded from this measurement as it is dependent on external factors (e.g., CPU, main memory, etc.). We measure the throughput in images/second as the average of 100 samples, and provide the 95% confidence intervals.
Our training procedure follows an identical curriculum to (yang2024kolmogorov), wherein we follow the hyperparameters of DeiT (touvron2021training). As such, we train using a batch size of 1024 with the AdamW optimizer (loshchilov2017decoupled). For the models, we apply Mimetic initialization to the attention layers (trockman2023mimetic) and set the patch size to 16. In GR-KAN, each of the 8 groups shares its 6 numerator coefficients and has 4 unique denominator coefficients. The first layer of GR-KAN has its group-wise rational function initialized to the identity function, and the second layer is initialized to a Swish function (ramachandran2017searching). For data augmentation and regularization techniques, we apply RandAugment (cubuk2020randaugment), Mixup (zhang2017mixup), CutMix (yun2019cutmix), Random Erasing (zhong2020random), weight decay, Label Smoothing (szegedy2016rethinking) and Stochastic Depth (huang2016deep). We conduct all experiments using an H200 GPU. The Appendix provides the complete set of model configurations and training hyperparameters.
@@
Reduced Gradient Rounding Error
@@
In addition to a significant speedup, the proposed FlashKAT also offers reduced rounding errors when computing the gradients for the group-wise rational function coefficients. This is a byproduct of our method, accumulating gradients in each block through a reduction sum rather than atomic adds. Table 5 demonstrates that the mean absolute error in gradient outputs in FlashKAT is orders of magnitude smaller than KAT. As such, FlashKAT is less susceptible to rounding errors that may contribute toward unstable gradient updates (e.g. low-precision training). Due to space limitations, we provide more details in the Appendix.
@@ 6 Conclusion
The successes of the KAN, due to its increased expressivity and interpretability, have made it a popular alternative to the MLP. However, it has yet to achieve widespread application in large-scale NLP and computer vision tasks due to its increased parameters, computational cost, lack of GPU hardware optimization, and training instability. The KAT has resolved many of the KAN’s shortcomings with the introduction of GR-KAN, but it still suffers from slow training speeds. In this work, we conduct experiments to understand the source of the slowdown, and we identifiy that the slow training speeds are a byproduct of memory stalls in the backward pass, a discovery that eluded previous works (yang2024kolmogorov; molina2019pad).
To address this bottleneck, we propose a novel restructuring of the group-wise rational function that allows efficient gradient accumulation by minimizing atomic add usage. The resulting FlashKAT achieves an 86.5x speedup compared to KAT, while reducing gradient rounding errors. The application of our work both enables wider adoption and expedites research on the KAT.
@@ Appendix A Appendix
Image Recognition Training Hyperparameters
@@
Model Layers Hidden size MLP size Heads Params
@@
Adam $\epsilon$ $\times 10^{-8} 1×1$$\times 10^{-8}$ $1\times 10^{-8}$
@@
Learning rate $\times 10^{-3} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$
@@
Reduced Gradient Rounding Error Expanded
FlashKAT also reduces rounding error when computing the gradients for the learned group-wise rational function coefficients, a byproduct of our implementation accumulating gradients in each block through sum reduction rather than using atomic adds. We validate that FlashKAT reduces gradient rounding errors by constructing an experiment measuring the coefficient gradient rounding errors. Our experiment generates $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, and $\mathbf{B}\in\mathbb{R}^{8\times 4}$ whose entries are distributed according to $\mathcal{N}(0,1)$. Then we compute $\mathbf{dA}\in\mathbb{R}^{8\times 6}$ and $\mathbf{dB}\in\mathbb{R}^{8\times 4}$ using the KAT group-wise rational function backward pass method in float64 and float32, and the FlashKAT group-wise rational function backward pass method in float32. We measure the mean absolute error (MAE) over 100 passes between the float64 gradient output and two float32 outputs. All these experiments were done on an RTX 4060 Ti. We report the mean, 95% confidence interval, and variance in Table 8.